### v1

@@ 1 Introduction
The traditional approach to dealing with finite strains in phenomenological elastoplasticity is based on the multiplicative decomposition of the deformation gradient [5]. The most common examples of adopting such an approach for modeling SMA mechanics include [6, 7, 8, 9, 10, 11, 12, 13]. The evolution of the elastic part is often governed by various isotropic hyperelastic potentials (e.g. Saint-Venant-Kirchhoff, Neo-Hookean); the remaining term(s) of the decomposition may vary in interpretation and they are related to, e.g., transformation, reorientation, conventional plasticity, viscoplasticity, or non-dissipative nonlinear deformation. Certain models incorporate the thermal component in the decomposition, thereby allowing for the simulation of thermomechanical coupling [14, 15, 16].
@@
The approach has been successfully applied to many constitutive models of metal plasticity [23] and damage [24, 25], polymers [26], wood [27], etc., and its limitations have also been addressed [28]. In the realm of SMA models, the methodology has been only recently applied in [29] and [30] for extending small-strain models [31] and [32] to finite strains, respectively.
@@ 2.2 Formulation within the Logarithmic Strain Space
[…]
@@
[…]
@@ 3.1 Finite Element Implementation
The logarithmic strain model of SMA has been implemented into finite element method in MATLAB computational environment. The backbone of the implementation is the lightweight code developed in [57], which features a time-efficient vectorized implementation. Its structure was modularized, and the elastoplastic core has been replaced by the present finite-strain SMA model. Moreover, the robustness has been reinforced by implementing a simple adaptive time-incrementation scheme.
[…]
@@ 3.4 Simulation 2: Stretching of a Helical Spring Coil
Computational analysis of the deformation of a helical spring is a common benchmark of finite-strain SMA models [11, 12, 13, 20]. We simulate the stretching of a single coil of a helical spring at a constant temperature (corresponding to the superelastic curve in Fig. 1), imposing periodic boundary conditions with a shifted axial displacement to ensure that the geometry represents one segment of an infinitely long virtual spring. The coil is wound from a wire with a circular cross-section of 0.1 mm in diameter. The outer diameter of the coil is 2.1 mm (i.e., the spring index 20), and the initial pitch is 0.1 mm (i.e., the initial pitch angle $\,{}^{\circ}$). The finite element model comprises 4,320 linear hexahedral elements with 48 elements per cross-section distributed uniformly in 90 layers along the wire axis. Stretching is simulated with a gradual change of prescribed displacement in the coil axis-direction of the central node at one end, while displacements in the other directions are fixed. The central node at the other end is fixed completely. Periodic boundary conditions with a shifted axial displacement are imposed at the remaining nodes of both ends (cross-sections) of the coil. The rigid body motion of the whole coil is prevented. Hence, the simulated response is the same in any cross-section as required by the infinite spring assumption.
@@ 4 Conclusions
We have extended a small-strain model based on [33] to the finite-strain regime using the consistent modular approach of [22]. The approach offers an alternative to both the “classical" multiplicative decomposition-based approach, which necessitates the complete rebuilding of the model structure and comprehensive mathematical manipulations, and to the logarithmic stress rate approach, which builds upon the original model but requires substantial computational effort linked with logarithmic co-rotational integration.
The logarithmic strain approach employed in this work preserves the original model structure, redirecting the modeling effort to the pre- and post-processing phases, with spectral decomposition of tensors being the primary computational burden. The modular structure of the logarithmic strain approach is also extremely beneficial for future modifications/development, e.g., towards capturing the plastic processes in NiTi martensite, cf. [64], or their coupling with phase transformation [45]. The first attempts in this direction have been performed in [65]. It is worth noting that the relative simplicity of implementation is counterbalanced by limited applicability in particularly severe loading modes; see [28] for details.

### v2

@@ 1 Introduction
The traditional approach to dealing with finite strains in phenomenological elastoplasticity is based on the multiplicative decomposition of the deformation gradient [5]. The most common examples of adopting such an approach for modeling SMA mechanics include [6, 7, 8, 9, 10, 11, 12]. The evolution of the elastic part is often governed by various isotropic hyperelastic potentials (e.g. Saint-Venant-Kirchhoff, Neo-Hookean); the remaining term(s) of the decomposition may vary in interpretation and they are related to, e.g., transformation, reorientation, conventional plasticity, viscoplasticity, or non-dissipative non-linear deformation. Certain models incorporate the thermal component in the decomposition, thereby allowing for the simulation of thermomechanical coupling [13, 14, 15].
@@
The approach has been successfully applied to many constitutive models of metal plasticity [22] and damage [23, 24], polymers [25], etc., and its limitations have also been addressed [26]. In the realm of SMA models, the methodology has been only recently applied in [27] and [28] for extending small-strain models [29] and [30] to finite strains, respectively.
@@ 2.2 Formulation within the Logarithmic Strain Space
where $k$ is a material parameter representing the maximum transformation strain of martensite attainable in uniaxial tensile loading and the material function $\langle\cdot\rangle$ takes a specific form proposed in [31]:
@@
$\Delta s^{AM}$ denotes the specific entropy change associated with the phase transformation, $T_{0}$ is the corresponding phase-equilibrium temperature (see [31] for details), and $\|\cdot\|$ denotes the Frobenius norm. Temperature, $T$, is considered a prescribed (albeit variable) parameter, as the superscript indicates.
@@ 3.1 Finite Element Implementation
The logarithmic strain model of SMA has been implemented into the finite element method in the MATLAB computational environment. The implementation builds on a lightweight open-source finite element code developed in [55]. The original code: i) features the innovative assembly of FEM matrices (discretized versions of operators), their memory-saving storage and efficient handling via full vectorization, ii) has a modular structure, which allows to alter the employed constitutive routine (state update procedure) in the same fashion as user material routines are provided in common commercial finite element packages (e.g. UMAT in Abaqus, USERMAT in Ansys, HYPELA2 in MSC MARC), and iii) is continuously extended, improved, and streamlined, cf. recent papers [56, 57, 58].
The fundamental structure of the code – boundary value problems are resolved via the finite element method with implicit time discretization and (semismooth) Newton solution algorithm for the linearized system – was retained. However, the following major modifications had to be made with respect to the finite-strain SMA model: 1) the elastoplastic constitutive routine (state update procedure) was substituted by the finite-strain SMA model described in Subsection 2.2, 2) the pre- and post-processing phases introduced in Subsections 2.1 and 2.3 were added to the main routine of the code; this required the substitution of the strain-displacement operator (matrix) by the deformation gradient-displacement operator and the construction of the projection tensors expressed by Eqs. (16) and (17), 3) the robustness of the code was increased by implementing an adaptive time-incrementation procedure, 4) the flexibility of the code was enhanced by the option to prescribe temperature within the integration (Gauss) points; this is particularly beneficial in SMA (isothermal) simulations, where the temperature is an additional, user-controlled parameter that fundamentally affects the mechanical response.
The main modification from the point of view of SMA modeling is the complete substitution of the constitutive routine listed above as 1). In their original work [55], the authors formulated two elastoplastic constitutive models using the conventional flow plasticity theory. Their state update procedure relied on the elastic predictor–plastic corrector (return mapping) solution algorithm. The proposed SMA model builds on works [31, 32] using GSM and its time-discrete formulation takes the compact form of the minimization problem presented in Eq. (15). Although this problem can be transformed into a system solvable with return mapping techniques, the minimization formulation naturally leads to a state update procedure based on (non-smooth) mathematical optimization. In the present study, the Matlab in-built function fminsearch based on the Nelder-Mead minimization algorithm is employed; this algorithm appeared robust and reliable in our previous studies. In future work, we plan to assess the performance of other minimization methods, e.g., bundle or non-smooth Newton methods. A Matlab-coded toolbox for automatic differentiation of mathematical functions called ADiGator was used to compute the consistent tangent modulus [59]. The robustness of the Newton method was reinforced by applying some regularization to functions $\langle\cdot\rangle$ and $d$ (defined in Eqs. (6) and (11)–(12), respectively). All numerical aspects of modifications of the original code stemming from its finite-strain extension will be detailed in a forthcoming paper [60].
@@ 3.4 Simulation 2: Stretching of a Helical Spring Coil
[…]
@@ 4 Conclusions
We have extended a small-strain model based on [31] to the finite-strain regime using the consistent modular approach of [21]. The approach offers an alternative to both the “classical" multiplicative decomposition-based approach, which necessitates the complete rebuilding of the model structure and comprehensive mathematical manipulations, and to the logarithmic stress rate approach, which builds upon the original model but requires substantial computational effort linked with logarithmic co-rotational integration. The logarithmic strain approach employed in this work preserves the original model structure, redirecting the modeling effort to the pre- and post-processing phases, with spectral decomposition of tensors being the primary computational burden. The modular structure of the logarithmic strain approach is also extremely beneficial for future modifications/development, e.g., towards capturing the plastic processes in NiTi martensite, cf. [66], or their coupling with phase transformation [43]. It is worth noting that the relative simplicity of implementation is counterbalanced by limited applicability in particularly severe loading modes; see [26] for details.
@@ Acknowledgements
[…]
Data Availability Data can be found at: https://doi.org/10.5281/zenodo.15396379.